OpenStudy (pagen13):

find the coefficient of the fourth term (x+2)^5

OpenStudy (anonymous):

okay you have to get x alone so you do the inverse operation to whatevers on the other side of x so you would have to do 5-2

OpenStudy (anonymous):

which is 3 and 3 is ur answer

OpenStudy (anonymous):

@Pagen13 did u understand that? :)

OpenStudy (pagen13):

3 isnt a choice..

OpenStudy (pagen13):

@ash_heather1

OpenStudy (anonymous):

um then what are ur choices?

OpenStudy (pagen13):

a.32 b.48 c.64 d.80

OpenStudy (anonymous):

um i dunno the answer then sry @JoeDeWise

OpenStudy (anonymous):

Okay, hang on.

OpenStudy (pagen13):

Okay...

OpenStudy (anonymous):

(x+2)5 =(x+2)5 =(x+2)*(x+2)*(x+2)*(x+2)*(x+2) =x5+10x4+40x3+80x2+80x+32

OpenStudy (pagen13):

so 32?

OpenStudy (anonymous):

yes. Pretty sure.

OpenStudy (pagen13):

Can you help with 3 more?

OpenStudy (anonymous):

I will try.

OpenStudy (pagen13):

find the coefficient of the third term of (3x-1)^5 a.248 b.270 c.360 d.505

OpenStudy (anonymous):

(3x−1)5 =(3x−1)5 =(3x−1)*(3x−1)*(3x−1)*(3x−1)*(3x−1) =243x5−405x4+270x3−90x2+15x−1

OpenStudy (anonymous):

B.

OpenStudy (pagen13):

Find the coefficient of the third term of (a+5b^2)^4 a.150 b.500 c.625 d.1,500

OpenStudy (anonymous):

(a+5b2)4 =625b8+500ab6+150a2b4+20a3b2+a4

OpenStudy (anonymous):

A

OpenStudy (anonymous):

?

OpenStudy (pagen13):

Sorry i had a phone call

OpenStudy (pagen13):

I have one more question but I am going to close this one a tag you in the next question

OpenStudy (anonymous):

okay tag me.